英語 (English)

基盤英語 (Basic English) (工 ((生 B)1 年))

マーシェソジェラルド · 准教授/ 鳴門教育大学 1 単位前期金 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】 Approximately half of each class will focus on Medical English.The

other half will focus on general conversational English. 【授業の概要】English for specific purposes (medical/dental English), General

conversational English 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 The purpose of this class is to improve students’ basic English skills.

The syllabus will be task based. Tasks will be integrated, targeting a variety of discrete skills. Grammar and vocabulary items will be addressed, but they will be viewed as subservient to meaning, rather than generative and devoid of context. The communicative aspect of language will be at the forefront.

【授業の計画】 1. Introductions / Getting to know your fellow students and teacher

2. Discussing language learning strategies and goals

3. Pronunciation workshop

4. How to take point form notes effectively

5. Using past tense to talk about events

6. Language for cause and effect, drawing conclusions

7. Expressing cause and effect in writing, developing a written argument

8. Preparation for presentations

9. Presentations

10. Using adjectives and descriptive language to describe cities and places

11. Describing your hometown, describing other cities

12. Pragmatically correct usage of various past tenses

13. Describing a vacation you have taken

14. Test practice

15. Test

16. Final class 【教科書】 Students will use material prepared by the instructor and self-generated

material. There will be no textbook. 【参考書等】 [参考資料] 【成績評価の方法】 Presentation: 20. Two written tests (about medical English): 20%

each. Speaking test: 20. Attendance and participation: 20 【再試験の有無】 Yes, if you have attended at least 2/3 of the class and your score is

at least 50% 【受講者へのメッセージ】 I hope that you find the medical English part of this class useful for your future, but you’ll have to study a lot of new vocabulary for the tests. For the general part of this class, you’ll have a chance to participate more actively and speak English as you complete various theme-based tasks.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220867 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ マーシェソ .

English

Basic English (工 ((生 B)1 年))

Gerard Marchesseau · Associate Professor / Naruto University of Education 1 unit 前期金 5 ・ 6

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ Approximately half of each class will focus on Medical English.The other

half will focus on general conversational English. Outline⟩ English for specific purposes (medical/dental English), General conver-

sational English �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ The purpose of this class is to improve students’ basic English skills. The

syllabus will be task based. Tasks will be integrated, targeting a variety of discrete skills. Grammar and vocabulary items will be addressed, but they will be viewed as subservient to meaning, rather than generative and devoid of context. The communicative aspect of language will be at the forefront.

Schedule⟩

1. Introductions / Getting to know your fellow students and teacher

2. Discussing language learning strategies and goals

3. Pronunciation workshop

4. How to take point form notes effectively

5. Using past tense to talk about events

6. Language for cause and effect, drawing conclusions

7. Expressing cause and effect in writing, developing a written argument

8. Preparation for presentations

9. Presentations

10. Using adjectives and descriptive language to describe cities and places

11. Describing your hometown, describing other cities

12. Pragmatically correct usage of various past tenses

13. Describing a vacation you have taken

14. Test practice

15. Test

16. Final class Textbook⟩ Students will use material prepared by the instructor and self-generated

material. There will be no textbook. Reference⟩ [参考資料] Evaluation Criteria⟩ Presentation: 20. Two written tests (about medical English):

20% each. Speaking test: 20. Attendance and participation: 20 Re-evaluation⟩ Yes, if you have attended at least 2/3 of the class and your score

is at least 50% Message⟩ I hope that you find the medical English part of this class useful for your future, but you’ll have to study a lot of new vocabulary for the tests. For the general part of this class, you’ll have a chance to participate more actively and speak English as you complete various theme-based tasks.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220867 Contact (Office-Hour, Room, E-mail)⟩

⇒ Marchesseau .